# The Angular Momentum Penrose Inequality

### v3

@@ Abstract
We prove the Angular Momentum Penrose Inequality: for asymptotically flat, axisymmetric initial data $(M^{3},g,K)$ satisfying the dominant energy condition with vacuum in the exterior region, and containing an outermost strictly stable marginally outer trapped surface (MOTS) $\Sigma$ of area $A$ and Komar angular momentum $J$,
@@
with equality if and only if the data arises from a slice of the Kerr spacetime.
The proof introduces a four-stage Jang–conformal–AMO method: (1) solve an axisymmetric Jang equation with twist as a lower-order perturbation; (2) solve an angular-momentum-modified Lichnerowicz equation; (3) establish angular momentum conservation via Stokes’ theorem for closed 2-forms; (4) apply the Dain–Reiris sub-extremality bound. The key innovation is the AM-Hawking mass $m_{H,J}(t):=\sqrt{m_{H}^{2}(t)+4\pi J^{2}/A(t)}$, which is monotonically non-decreasing along the $p$-harmonic flow and converges to $M_{\mathrm{ADM}}$.
As a methodological application, we outline a new proof strategy for the Charged Penrose Inequality $M_{\mathrm{ADM}}\geq M_{\mathrm{irr}}+Q^{2}/(4M_{\mathrm{irr}})$ for non-rotating Einstein–Maxwell data, demonstrating the versatility of the Jang–conformal–AMO framework. This charged case is presented as a proof outline; complete rigorous proofs exist in the literature [37, 71].
Keywords: Penrose inequality, angular momentum, Kerr spacetime, marginally outer trapped surfaces, Jang equation, conformal method, positive mass theorem
@@ 1 Introduction
1.1 Historical Context and Physical Motivation
The Penrose inequality, conjectured by Roger Penrose in 1973 [48], encapsulates a fundamental principle of black hole physics: black holes cannot be “underweight” for their size. It relates the ADM mass of an asymptotically flat spacetime to the area of its black hole horizons:
@@ 1.1 Historical Context and Physical Motivation
where $A$ is the area of the outermost marginally outer trapped surface (MOTS). This inequality was established for time-symmetric (Riemannian) initial data by Huisken–Ilmanen [32] using inverse mean curvature flow and by Bray [10] using conformal flow. The spacetime (non-time-symmetric) case has been studied extensively using the Jang equation approach [11, 29].
However, the classical formulation (1) does not account for the angular momentum of the black hole. For rotating (Kerr) black holes, angular momentum directly affects the horizon structure and is a conserved quantity under Einstein evolution.
Definition 1.1 (Sub-Extremality).
A Kerr black hole with mass $M$ and angular momentum $J=aM$ is called sub-extremal if $|a|<M$, extremal if $|a|=M$, and super-extremal (or naked singularity) if $|a|>M$. Equivalently, in terms of the dimensionless spin $\chi:=a/M=J/M^{2}$: sub-extremal means $|\chi|<1$. For an axisymmetric MOTS with area $A$ and Komar angular momentum $J$, the sub-extremality condition is $A\geq 8\pi|J|$, which is equivalent to the existence of a Kerr solution with matching $(A,J)$. The sub-extremality factor appearing in monotonicity formulas is:
$1-\frac{64\pi^{2}J^{2}}{A^{2}}=1-\left(\frac{8\pi|J|}{A}\right)^{2}.$
Key algebraic fact: This factor is non-negative if and only if $A\geq 8\pi|J|$. Indeed, $1-(8\pi|J|/A)^{2}\geq 0$ is equivalent to $(8\pi|J|/A)^{2}\leq 1$, i.e., $8\pi|J|\leq A$. At the extremal limit ($A=8\pi|J|$), the factor vanishes; for sub-extremal configurations ($A>8\pi|J|$), it is strictly positive. The Dain–Reiris inequality [22] ensures this factor is non-negative for all stable MOTS in axisymmetric data satisfying DEC.
The Kerr solution with mass $M$ and angular momentum $J=aM$ (where $a$ is the spin parameter with $|a|\leq M$ for sub-extremal black holes; see Definition 1.1) has horizon area
@@
which depends nontrivially on the spin parameter $a$. This motivates the search for a generalized Penrose inequality that incorporates both horizon area and angular momentum.
@@ 1.2 Main Result
We prove the natural extension incorporating angular momentum:
@@
    Dominant energy condition: $\mu\geq|\boldsymbol{j}|_{g}$, where
@@
    is the energy density and $\boldsymbol{j}$ is the momentum density vector field (see Remark 1.9);
@@
    Axisymmetry: There exists a Killing field $\eta=\partial_{\phi}$ generating rotations, with $\eta\neq 0$ on $M\setminus\Gamma$ where $\Gamma$ denotes the rotation axis; [^1] The axis $\Gamma=\{\eta=0\}$ is a 1-dimensional submanifold (possibly with multiple components) where the Killing field vanishes. The condition $\eta\neq 0$ on $M\setminus\Gamma$ ensures the orbits of $\eta$ are circles, corresponding to physical rotation about the axis.
@@
    Vacuum in exterior region: The constraint equations hold with $\mu=|\boldsymbol{j}|=0$ in the exterior region $M_{\mathrm{ext}}:=M\setminus\overline{\mathrm{Int}(\Sigma)}$, where $\mathrm{Int}(\Sigma)$ denotes the bounded component of $M\setminus\Sigma$. This hypothesis is essential for angular momentum conservation along the flow (see Remark 1.13);
@@
    Strictly stable outermost MOTS: There exists an outermost MOTS $\Sigma\subset M$ that is strictly stable, i.e., the principal eigenvalue of the MOTS stability operator (Definition 4.4) satisfies $\lambda_{1}(L_{\Sigma})>0$.
Let $A:=\int_{\Sigma}dA_{g}$ denote the area of $\Sigma$ with respect to the physical metric $g$. Let $\nu$ denote the outward-pointing unit normal to $\Sigma$ (i.e., pointing toward spatial infinity, satisfying $\langle\nu,\nabla r\rangle>0$ asymptotically for any radial coordinate $r$). Define the Komar angular momentum:
$J:=\frac{1}{8\pi}\int_{\Sigma}K(\eta,\nu)\,d\sigma.$
This orientation convention ensures $J>0$ for prograde rotation (angular momentum aligned with the positive $\phi$-direction). The Komar definition agrees with the ADM angular momentum at infinity for axisymmetric asymptotically flat data with decay rate $\tau>1/2$ (Definition 4.2); see [16, 37] for the equivalence under these decay conditions.
Then:
@@
with equality if and only if the initial data arises from a slice of the Kerr spacetime with parameters $(M,a=J/M)$. The equality characterization is proven in Theorem 9.1 (Section 9) using the Mars–Simon tensor.
Remark 1.3 (Scope of the Result).
Theorem 1.2 establishes the Angular Momentum Penrose Inequality (2) for initial data satisfying axisymmetry and a vacuum exterior. The axisymmetry is required for defining the Komar angular momentum and for the orbit-space reduction of the Jang equation. The vacuum condition ensures $d^{\dagger}\alpha_{J}=0$, which is essential for angular momentum conservation (Theorem 6.14). We discuss these restrictions further in Remark 1.6 and Section 10.
@@
The proof relies on several established results: the Jang equation existence theory of Han–Khuri [29], the area-angular momentum inequality of Dain–Reiris [22], the $p$-harmonic flow monotonicity of Agostiniani–Mazzieri–Oronzio [1], and the Mars–Simon characterization of Kerr [83, 84]. Our contribution is the synthesis of these methods and the introduction of the AM-Hawking mass functional.
Remark 1.5 (Role of Each Hypothesis).
The hypotheses (H1)–(H4) enter the proof at specific points:
[…]
  • (H2) Axisymmetry: Essential for defining Komar angular momentum and for the orbit-space reduction of the Jang equation (Theorem 4.12). Also enables the twist perturbation analysis (Lemma 4.14).
  • (H3) Exterior vacuum: Critical for angular momentum conservation along the AMO flow (Theorem 6.14). Without vacuum, there would be matter fluxes that could change $J$.
  • (H4) Strictly stable MOTS: Used in Stage 1 to construct the Jang solution with controlled logarithmic blow-up and cylindrical ends (Theorem 4.12). The spectral gap $\lambda_{1}(L_{\Sigma})>0$ ensures Fredholm theory applies.
Remark 1.6 (Scope and Limitations).
The result is presently restricted to axisymmetric data sets with a vacuum exterior. These two constraints are fundamental to the proof:
  1. 1.
    Vacuum requirement: The result is strictly limited to data with $\mu=|\boldsymbol{j}|=0$ in the exterior region. This is necessary for the conservation of $J$ along the flow (Theorem 6.14). In the presence of matter, the Komar angular momentum would drift, and the inequality might require modification—see Remark 1.13 for details.
  2. 2.
    Axisymmetry requirement: The proof relies on the existence of the Killing field $\eta=\partial_{\phi}$, without which the Komar angular momentum $J$ is undefined. The non-axisymmetric case remains a major open problem: there is no canonical definition of quasi-local angular momentum, and the twist perturbation analysis does not apply.
Dynamical horizons and the case of multiple black holes are discussed as open problems in Section 10.
Physical situations excluded by the hypotheses:
  • Matter outside the horizon: Accretion disks, neutron star matter, or any stress-energy in the exterior region violates (H3). Angular momentum transport by matter would invalidate $J$-conservation.
  • Multiple disconnected horizons: The theorem addresses a single outermost MOTS. Binary black hole configurations require the “area additivity” conjecture (see Section 10).
  • Non-axisymmetric perturbations: Gravitational wave perturbations, tidally deformed horizons, or any configuration lacking a rotational Killing field violates (H2). The quasi-local angular momentum is then undefined.
  • Marginally stable MOTS ($\lambda_{1}=0$): The strictly stable condition (H4) excludes extremal or marginally stable horizons. The function space analysis requires $\lambda_{1}>0$ for Fredholm theory.
  • Dynamical horizons: Time-dependent horizons (e.g., during merger) are not covered; the theorem applies to a fixed initial data slice.
  • Charged rotating black holes (Kerr-Newman): The combined case $J\neq 0$ and $Q\neq 0$ is conjectured (Conjecture 10.36) but not proven.
Corollary 1.7 (Quantitative Deficit Bound).
Under the hypotheses of Theorem 1.2, define the AM-Penrose deficit:
@@
Then:
  1. (i)
    Lower bound in terms of Kerr deviation: If $\mathcal{S}_{(g,K)}\neq 0$ (the Kerr deviation tensor from Definition 1.11), then
    $\delta_{PI}\geq c_{0}\int_{M}|\mathcal{S}_{(g,K)}|^{2}\,dV_{g}$
    for an explicit constant $c_{0}>0$ depending on the geometry. Note: this bound involves the Kerr deviation tensor, not the raw $\sigma^{TT}$, since Kerr slices themselves have $\sigma^{TT}\neq 0$.
  2. (ii)
    Rigidity: $\delta_{PI}=0$ if and only if $(M,g,K)$ is isometric to a slice of Kerr with $(M,a=J/M)$.
  3. (iii)
    Stability bound: For data $(g_{\epsilon},K_{\epsilon})$ that is $C^{2}$-close to Kerr with parameters $(M,a)$:
    $\left|M_{\mathrm{ADM}}(g_{\epsilon})-\sqrt{\frac{A_{\epsilon}}{16\pi}+\frac{4\pi J_{\epsilon}^{2}}{A_{\epsilon}}}\right|\leq C\|(g_{\epsilon},K_{\epsilon})-(g_{\mathrm{Kerr}},K_{\mathrm{Kerr}})\|_{C^{2}}$
    for an explicit constant $C$ depending on $(M,a)$.
Proof sketch.
Part (i) follows from the rigidity analysis: $\delta_{PI}=0$ requires $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}=0$ identically (Section 9). The quantitative version comes from tracking the mass deficit through the Jang–conformal construction.
Part (ii) is proven in Theorem 9.1.
Part (iii) follows from the continuous dependence of ADM mass on the metric in appropriate norms, combined with the explicit Kerr calculation (Theorem 2.3). ∎
Remark 1.8 (Regularity Requirements).
Theorem 1.2 requires the following regularity:
  1. (i)
    Metric and extrinsic curvature: $(g,K)\in C^{k,\beta}_{\mathrm{loc}}(M)\times C^{k-1,\beta}_{\mathrm{loc}}(M)$ for some $k\geq 3$ and $\beta\in(0,1)$. This ensures:
    • Well-definedness of scalar curvature $R_{g}\in C^{k-2,\beta}$;
    • Elliptic regularity for the Jang equation (Theorem 4.12);
    • $C^{1,\beta}$ regularity of $p$-harmonic potentials via Tolksdorf–Lieberman theory.
  2. (ii)
    Asymptotic flatness: The decay conditions in Definition 4.2 with $\tau>1/2$ and $k\geq 3$ ensure well-defined ADM mass.
  3. (iii)
    MOTS regularity: The outermost MOTS $\Sigma$ is a $C^{k,\beta}$ embedded surface (automatic from elliptic regularity when $g\in C^{k,\beta}$).
  4. (iv)
    Minimal regularity: The proof can be extended to $C^{2}$ metrics using distributional techniques, but we state Theorem 1.2 for $C^{3,\beta}$ data for clarity.
The Lockhart–McOwen theory for weighted Sobolev spaces (Definition 5.1) provides the precise functional-analytic framework.
Remark 1.9 (Notation: Angular Momentum vs. Momentum Density).
We use two distinct quantities with visually distinct notation to avoid confusion:
  • $J$ (roman, scalar): The Komar angular momentum, defined as the surface integral $J=\frac{1}{8\pi}\int_{\Sigma}K(\eta,\nu)\,d\sigma$. This is the total angular momentum of the black hole.
  • $\boldsymbol{j}$ (boldface, vector field): The momentum density from the constraint equations, defined by $\boldsymbol{j}_{i}=D^{k}K_{ki}-D_{i}(\mathrm{tr}K)$. Its norm $|\boldsymbol{j}|_{g}$ appears in the dominant energy condition.
For vacuum data, $\boldsymbol{j}=0$ identically, so the DEC reduces to $\mu\geq 0$.
Additional notation clarifications:
  • $\beta$ (in $C^{k,\beta}$): The Hölder exponent, a regularity parameter $\beta\in(0,1)$ appearing in function space definitions. We use $\beta$ rather than the traditional $\alpha$ to avoid any confusion with the Komar 1-form.
  • $\alpha_{J}$: The Komar 1-form, defined as $\alpha_{J}=\frac{1}{8\pi}K(\eta,\cdot)^{\flat}_{g}$. Its integral over a surface gives the angular momentum: $J=\int_{\Sigma}\star_{g}\alpha_{J}$.
This notational convention eliminates any potential ambiguity between regularity exponents and angular momentum-related quantities.
Remark 1.10 (Essential Role of Each Hypothesis).
  • (H1) DEC ensures $R_{\bar{g}}\geq 0$ on the Jang manifold via the Bray–Khuri identity.
  • (H2) Axisymmetry enables the definition of Komar angular momentum and ensures the AMO flow preserves the symmetry.
  • (H3) Vacuum is critical: it ensures the Komar form is co-closed ($d^{\dagger}\alpha_{J}=0$), which implies $d(\star\alpha_{J})=0$ and hence angular momentum conservation (Theorem 6.14).
  • (H4) Stability ensures the Jang equation has the correct blow-up behavior and the Dain–Reiris inequality $A\geq 8\pi|J|$ holds.
@@
A common misconception is that $\sigma^{TT}=0$ characterizes Kerr. This is false:
  • Kerr slices have $\sigma^{TT}\neq 0$: Boyer–Lindquist slices of Kerr are not conformally flat. The induced 3-metric has non-trivial Cotton tensor, and the extrinsic curvature has genuine TT-content encoding frame-dragging.
  • Bowen–York data has $\sigma^{TT}=0$: Bowen–York initial data [89] is conformally flat with $\sigma^{TT}=0$, but it does not represent a Kerr slice—its evolution produces gravitational radiation.
The correct characterization uses the Mars–Simon tensor, which vanishes for Kerr (any slice) but is non-zero for Bowen–York and other non-Kerr configurations.
@@
Relaxing the vacuum hypothesis to DEC-only for $J\neq 0$ would require either:
  1. (a)
    A modified monotonicity formula that tracks $J(t)$ variations and bounds their contribution—this appears technically challenging as no candidate formula is known.
  2. (b)
    Restricting to matter models with $\boldsymbol{j}_{\phi}=0$, e.g., perfect fluids co-rotating with the symmetry. This is a non-trivial physical assumption beyond DEC.
We therefore view vacuum as the minimal natural hypothesis for the angular momentum Penrose inequality in the present framework. Without vacuum, the Komar angular momentum $J$ is not conserved along homologous surfaces, rendering the inequality $M\geq f(A,J)$ ill-posed: which value of $J$ (horizon vs. ADM vs. intermediate) should appear? The charged extension (§10.1) shows how specific matter models (electrovacuum) can be incorporated when their angular momentum contributions are computable.
@@
The vacuum exterior hypothesis (H3) is physically reasonable for isolated black holes in astrophysical settings:
  1. 1.
    Event horizon vicinity: In the region immediately outside a stationary black hole, matter cannot remain in equilibrium without extraordinary support—it either falls into the black hole or is ejected. The “vacuum zone” near the horizon is therefore a generic feature of isolated black holes.
  2. 2.
    Astrophysical black holes: Real astrophysical black holes (e.g., Sgr A*, M87*) are surrounded by accretion disks, but the matter density falls off rapidly with distance from the disk midplane. The region swept by the AMO flow can be chosen to avoid dense matter concentrations.
  3. 3.
    Gravitational wave events: In binary black hole mergers (LIGO/Virgo observations), the pre-merger spacetime is vacuum outside the individual horizons. The inequality applies to initial data representing snapshots of such systems.
  4. 4.
    Cosmic censorship context: The Penrose inequality is fundamentally a statement about gravitational collapse leading to black hole formation. In such scenarios, matter has already collapsed into the singularity; the exterior region is vacuum by the time a stable horizon forms.
@@
  1. (1)
    Squared form:
    $M_{\mathrm{ADM}}^{2}\geq\frac{A}{16\pi}+\frac{4\pi J^{2}}{A}$
    Obtained by squaring (2). This form is often more convenient for computations.
  2. (2)
    Irreducible mass form: With $M_{irr}=\sqrt{A/(16\pi)}$:
    $M_{\mathrm{ADM}}^{2}\geq M_{irr}^{2}+\frac{J^{2}}{4M_{irr}^{2}}$
    This form emphasizes the decomposition into irreducible mass and rotational contribution.
  3. (3)
    Area bound form: Rearranging gives the area lower bound
@@
    when $|J|\leq M_{\mathrm{ADM}}^{2}$ (sub-extremality). This matches $A_{\text{Kerr}}(M,a)$ with $a=J/M$.
Validity: All three forms are equivalent for any configuration satisfying the theorem’s hypotheses. The sub-extremality condition $|J|\leq M_{\mathrm{ADM}}^{2}$ required for form (3) is automatically satisfied for physical black holes by the Dain–Reiris inequality $A\geq 8\pi|J|$ combined with the Penrose inequality—see Theorem 7.2.
@@
When $J=0$ (time-symmetric or non-rotating data), Theorem 1.2 reduces to the standard Penrose inequality (1):
$M_{\mathrm{ADM}}\geq\sqrt{\frac{A}{16\pi}+0}=\sqrt{\frac{A}{16\pi}}.$
This includes:
  • Time-symmetric data ($K=0$): Here $J=0$ trivially, and Theorem 1.2 reproduces the Riemannian Penrose inequality proved by Huisken–Ilmanen [32] and Bray [10].
  • Axisymmetric data with vanishing twist: Even with $K\neq 0$, if the twist $\omega_{ij}=K_{i\phi}\delta_{j}^{\phi}-K_{j\phi}\delta_{i}^{\phi}$ vanishes or integrates to zero over $\Sigma$, the Komar integral gives $J=0$.
  • Spherically symmetric data: Spherical symmetry implies $J=0$ by parity, so Theorem 1.2 gives the Schwarzschild bound.
The condition $J=0$ simplifies the proof significantly: Stage 3 (angular momentum conservation) becomes trivial, and the monotonicity reduces to the standard Hawking mass monotonicity. Our proof is thus consistent with and generalizes existing results.
1.3 Significance and Relation to Prior Work
Theorem 1.2 establishes a geometric inequality incorporating both horizon area and angular momentum for axisymmetric vacuum initial data. We describe the contributions of this paper and their relationship to prior work, with appropriate context.

### v4

@@ Abstract
We establish the Penrose inequality with angular momentum for asymptotically flat, axisymmetric vacuum initial data containing a strictly stable marginally outer trapped surface. Specifically, if $(M^{3},g,K)$ satisfies the dominant energy condition and $\Sigma$ denotes the outermost MOTS with area $A$ and Komar angular momentum $J$, then
@@
with equality characterizing slices of Kerr. The argument combines the Jang equation approach of Bray–Khuri and Han–Khuri with the $p$-harmonic level set method of Agostiniani–Mazzieri–Oronzio. A central role is played by a modified Hawking mass that incorporates angular momentum and remains monotone along the flow. The conservation of $J$ under this flow relies on the vacuum hypothesis through a cohomological argument. We also characterize the equality case using the Mars–Simon tensor.
Keywords: Penrose inequality, angular momentum, Kerr spacetime, MOTS, Jang equation, conformal method
@@ 1 Introduction
1.1 Background
The Penrose inequality, proposed by Penrose in 1973 [48], asserts that the ADM mass of an asymptotically flat spacetime bounds from below a quantity determined by its black hole horizons:
@@ 1.1 Background
where $A$ denotes the area of the outermost marginally outer trapped surface. This was established for time-symmetric data by Huisken–Ilmanen [32] using inverse mean curvature flow and by Bray [10] via conformal flow. The general (non-time-symmetric) case has been studied by Bray–Khuri [11] and Han–Khuri [29] using the Jang equation.
The bound (1) does not account for angular momentum. For rotating black holes, however, angular momentum affects the horizon geometry in an essential way. The Kerr solution with mass $M$ and angular momentum $J=aM$ has horizon area
@@
which depends on the spin parameter $a$. This suggests that the correct generalization should incorporate both area and angular momentum.
Recall that a Kerr black hole is sub-extremal when $|a|<M$ (equivalently, when the dimensionless spin $\chi=J/M^{2}$ satisfies $|\chi|<1$), extremal when $|a|=M$, and would be a naked singularity if $|a|>M$. For an axisymmetric MOTS, the analogous sub-extremality condition is $A\geq 8\pi|J|$, ensured by the Dain–Reiris inequality [22] for stable MOTS in axisymmetric data satisfying the dominant energy condition.
@@ 1.2 Main Result
Our main theorem establishes the natural generalization:
@@
    $\mu\geq|\boldsymbol{j}|_{g}$ (dominant energy condition), where
@@
    and $\boldsymbol{j}$ is the momentum density (see Remark 1.7);
@@
    There exists a Killing field $\eta=\partial_{\phi}$ generating rotations (axisymmetry);
@@
    $\mu=|\boldsymbol{j}|=0$ in the exterior region $M\setminus\overline{\mathrm{Int}(\Sigma)}$ (vacuum);
@@
    The outermost MOTS $\Sigma$ is strictly stable, i.e., $\lambda_{1}(L_{\Sigma})>0$.
Let $A$ be the area of $\Sigma$ and let
$J:=\frac{1}{8\pi}\int_{\Sigma}K(\eta,\nu)\,d\sigma$
be the Komar angular momentum, where $\nu$ is the outward unit normal. Then
@@
with equality if and only if the data arises from a slice of Kerr.
The vacuum hypothesis (H3) is used in an essential way: it ensures conservation of angular momentum along the flow (Theorem 6.14). Without this assumption, matter could transport angular momentum and the monotonicity argument would fail. We discuss this further in Remark 1.10.
The Komar angular momentum agrees with the ADM angular momentum at infinity for axisymmetric asymptotically flat data with decay rate $\tau>1/2$; see [16, 37]. The orientation convention is that $J>0$ for prograde rotation. The equality case is characterized in Theorem 9.1 using the Mars–Simon tensor.
@@
The proof builds on several established results: Han–Khuri’s Jang equation existence theory [29], the Dain–Reiris area-angular momentum inequality [22], the Agostiniani–Mazzieri–Oronzio $p$-harmonic monotonicity [1], and the Mars–Simon characterization of Kerr [83, 84]. Our contribution is the synthesis and the introduction of the angular-momentum-corrected Hawking mass.
Remark 1.3 (Role of the Hypotheses).
The dominant energy condition (H1) yields $R_{\bar{g}}\geq 0$ on the Jang manifold via the Bray–Khuri identity, which drives the monotonicity. Axisymmetry (H2) is needed to define $J$ and to reduce the Jang equation to an orbit-space problem. The vacuum condition (H3) ensures the Komar form is co-closed, giving angular momentum conservation. Strict stability (H4) controls the Jang blow-up rate and gives the spectral gap needed for Fredholm theory on the cylindrical end.
Remark 1.4 (Limitations).
The theorem requires axisymmetry and exterior vacuum. These are genuine restrictions: without vacuum, the Komar angular momentum can drift along the flow, since even data satisfying the dominant energy condition may have non-zero azimuthal momentum flux; without axisymmetry, there is no Killing field and quasi-local angular momentum is not canonically defined. The cases of multiple horizons, dynamical horizons, and non-axisymmetric data remain open; see Section 10.
Corollary 1.5 (Deficit Bound).
Under the hypotheses of Theorem 1.1, define
@@
Then $\delta_{PI}=0$ if and only if the data is Kerr. For data $C^{2}$-close to Kerr, the deficit is controlled by the deviation from Kerr geometry.
Remark 1.6 (Regularity).
We assume $(g,K)\in C^{k,\beta}_{\mathrm{loc}}\times C^{k-1,\beta}_{\mathrm{loc}}$ for $k\geq 3$ and $\beta\in(0,1)$, with asymptotic flatness as in Definition 4.2. This ensures elliptic regularity for the Jang and Lichnerowicz equations, and well-defined ADM mass.
Remark 1.7 (Notation).
To avoid confusion: $J$ (roman) is the Komar angular momentum, a scalar; $\boldsymbol{j}$ (boldface) is the momentum density vector field from the constraints. For vacuum data $\boldsymbol{j}=0$. We also use $\alpha_{J}=\frac{1}{8\pi}K(\eta,\cdot)^{\flat}$ for the Komar 1-form and $\beta\in(0,1)$ for Hölder exponents. We denote the normalized Willmore functional by $W(t)=\frac{1}{16\pi}\int_{\Sigma_{t}}H^{2}dA$, which is dimensionless and satisfies $W\leq 1$ for spherical surfaces in non-negative scalar curvature manifolds.
@@
A common misconception is that $\sigma^{TT}=0$ characterizes Kerr. This is false: Boyer–Lindquist slices of Kerr are not conformally flat, and the induced 3-metric has non-trivial Cotton tensor with extrinsic curvature encoding frame-dragging, so Kerr slices have $\sigma^{TT}\neq 0$. Conversely, Bowen–York initial data [89] is conformally flat with $\sigma^{TT}=0$, but it does not represent a Kerr slice—its evolution produces gravitational radiation. The correct characterization uses the Mars–Simon tensor, which vanishes for Kerr (any slice) but is non-zero for Bowen–York and other non-Kerr configurations.
@@
Relaxing the vacuum hypothesis to DEC-only for $J\neq 0$ would require either a modified monotonicity formula that tracks $J(t)$ variations and bounds their contribution (which appears technically challenging as no candidate formula is known), or restricting to matter models with $\boldsymbol{j}_{\phi}=0$, such as perfect fluids co-rotating with the symmetry (which is a non-trivial physical assumption beyond DEC). We therefore view vacuum as the minimal natural hypothesis for the angular momentum Penrose inequality in the present framework. Without vacuum, the Komar angular momentum $J$ is not conserved along homologous surfaces, rendering the inequality $M\geq f(A,J)$ ill-posed: which value of $J$ (horizon vs. ADM vs. intermediate) should appear? The charged extension (§10.1) shows how specific matter models (electrovacuum) can be incorporated when their angular momentum contributions are computable.
@@
The vacuum exterior hypothesis (H3) is physically reasonable for isolated black holes in astrophysical settings. In the region immediately outside a stationary black hole, matter cannot remain in equilibrium without extraordinary support—it either falls into the black hole or is ejected, making the “vacuum zone” near the horizon a generic feature. Real astrophysical black holes (e.g., Sgr A*, M87*) are surrounded by accretion disks, but the matter density falls off rapidly with distance from the disk midplane, so the region swept by the AMO flow can be chosen to avoid dense matter concentrations. In binary black hole mergers (LIGO/Virgo observations), the pre-merger spacetime is vacuum outside the individual horizons, and the inequality applies to initial data representing snapshots of such systems. Moreover, in the cosmic censorship context, the Penrose inequality is fundamentally a statement about gravitational collapse leading to black hole formation; in such scenarios, matter has already collapsed into the singularity, and the exterior region is vacuum by the time a stable horizon forms.
@@
The squared form $M_{\mathrm{ADM}}^{2}\geq A/(16\pi)+4\pi J^{2}/A$ is obtained by squaring (2) and is often more convenient for computations. The irreducible mass form, with $M_{irr}=\sqrt{A/(16\pi)}$, reads $M_{\mathrm{ADM}}^{2}\geq M_{irr}^{2}+J^{2}/(4M_{irr}^{2})$ and emphasizes the decomposition into irreducible mass and rotational contribution. The area bound form rearranges the inequality to give
@@
when $|J|\leq M_{\mathrm{ADM}}^{2}$ (sub-extremality); this matches $A_{\text{Kerr}}(M,a)$ with $a=J/M$.
All three forms are equivalent for any configuration satisfying the theorem’s hypotheses. The sub-extremality condition $|J|\leq M_{\mathrm{ADM}}^{2}$ required for the area bound form is automatically satisfied for physical black holes by the Dain–Reiris inequality $A\geq 8\pi|J|$ combined with the Penrose inequality—see Theorem 7.2.
@@
When $J=0$, the theorem reduces to the standard Penrose inequality $M_{\mathrm{ADM}}\geq\sqrt{A/(16\pi)}$, which includes time-symmetric data ($K=0$), axisymmetric data with vanishing twist, and spherically symmetric data. In these cases Stage 3 of the proof (angular momentum conservation) becomes trivial.